A sample of methane gas has a volume of 38mL at pressure of 2.2atm. What is the volume of gas in mL at 6.1 atm.
1
Expert's answer
2019-11-14T06:17:00-0500

From the Boyle–Mariotte law:

P1V1 = P2V2

V2 = P1×V1/V2

V2 = 2,2×38/6,1 = 13,7 ml